# A Comprehensive Survey on Automatic Text Summarization with Exploration of LLM-Based Methods

### v2

@@ 6.1.1 Extractive Summarization
Extractive ATS based on supervised learning uses a classifier to categorize sentences as summaries or non-summaries. The process involves representing sentences, classifying them, and selecting those labeled as summaries. Classifiers from other tasks can be adapted for this purpose, with options including machine-learning-based (ML) and deep-learning-based (DL) methods. While supervised approaches improve sentence representation and extraction accuracy using labeled data, they incur additional costs for annotation.
@@
DL-based Models: Deep learning models, particularly pre-trained ones, offer rich textual representations that enhance sentence vector construction and improve summarization accuracy. [197] introduced SqueezeBERTSum, fine-tuned with a SqueezeBERT encoder variant, while [198] explored generative and discriminative techniques to fuse domain knowledge into knowledge adapters, integrating them into pre-trained models for efficient fine-tuning. However, training these models is computationally intensive, requiring significant hardware resources.
@@
6.1.2 Abstractive Summarization
Abstractive summarization models generate summaries by producing sentences distinct from the original text. This can be achieved through text structuring and combining or via a generative models that being trained on predicting the next token. Abstractive summarization can be classified as either “rule-based” or “generative” methods. The The Rule-based Summarization methods will be introduced as the following:
@@ 6.1.3 Hybrid Summarization
Prompt Engineering Template Engineering [15, zhang-etal-2024-benchmarking, 247, 248, 16, 249] Predictable output, easy to control structure. Limited flexibility, may not capture complex relationships.
@@
Retrieval-Augmented Generation [223, 254, 249, 255, 256] Can use existing knowledge bases effectively. Reliability depends on retrieval accuracy; may introduce bias.
Fine-tuning Internal parameters [69, 257] Tailors model to specific tasks or domains. Destroying the original structure and capabilities of LLM.
External adapters [258, 259, 260] Less risk of overfitting; can adapt quickly to new tasks. May not transfer improvements as effectively as full fine-tuning.
@@
6.2 LLM-based Methods
@@
In ATS, LLMs have achieved results comparable to or exceeding human performance. Studies indicate that their summaries match human-crafted ones in quality and paraphrasing diversity [264, 265, zhang-etal-2024-benchmarking, 266, 267]. Moreover, when prompted with simple task descriptions, LLMs generate preferred and highly factual summaries, avoiding dataset-specific pitfalls [268]. This review bridges gaps in the literature by providing a comprehensive overview of LLMs in ATS.
OpenAI’s adaptation of GPT-3 using reinforcement learning marked a breakthrough in LLM-based summarization [269]. Subsequent research has further advanced the field [270]. LLM-based methods can be generally categorized as: 1) prompt engineering, 2) fine-tuning, and 3) knowledge distillation. The advantages and limitations of these methods are summarized in Table 6.
6.2.1 Prompt Engineering for LLM-based ATS
@@
$S=LLM_{\theta}\left(G(p)+x\right)$ (3)
where, $G$ refines the initial prompt $p$ into an optimized form through methods such as human design, supplementary retrieval, or generation by LLMs. The refined prompt is then combined with raw input text $x$ to leverage the knowledge encoded in the LLM, with $\theta$ representing the model’s parameters. This approach improves the interpretation of input text [15], resulting in higher-quality summaries $S$.
@@
The key advantage of prompt engineering is its efficiency, as it minimizes the need for extensive training and can operate effectively with a small set of examples [16]. This study examines various implementations of prompt engineering in ATS, focusing on techniques like Template Engineering, Chain of Thought (CoT), Agent Interactions, and Retrieval-Augmented Generation (RAG).
Template Engineering: The most straightforward approach to creating prompts is to manually design intuitive templates based on our human knowledge. Examples of such templates include:
@@
    Article: $[Input]$. Summarize the article in three sentences. Summary: [zhang-etal-2024-benchmarking]
@@
Chain of Thought: Chain of Thought (CoT) refers to a series of intermediate reasoning steps to allow LLMs handling complex reasoning tasks [272]. To address issues like factual hallucination and redundancy in ATS, [240] proposed the Summary Chain-of-Thought (SumCoT) technique, which guides LLMs to generate summaries step by step. This approach enables the integration of fine-grained details from source documents into the final summaries, mimicking the human writing process. Similarly, [250] introduced the ”Chain of Density” (CoD) method to create detailed, entity-centric summaries. Experiments showed that CoD-generated summaries are more abstractive, exhibit better fusion, and reduce lead bias compared to GPT-4 summaries generated with a standard prompt. Additionally, [235] explored tuning LLMs with a chain-of-thought strategy to produce summaries for individual comment clusters, advancing the application of CoT in ATS.
@@
Agent Interactions: Agents are artificial entities that sense their environment, make decisions, and take actions [273]. In ATS, [251] proposed a tri-agent pipeline (generator, instructor, editor) to customize LLM-generated summaries to better meet user expectations. [101] used agents to address dialogue summarization challenges like lengthy inputs, content verification, limited annotated data, and evaluation, with GPT-3 serving as an offline data annotator to handle dataset scarcity and privacy constraints. [252] employed clustering-based agents to improve intent detection by abstracting key elements while removing irrelevant information. Similarly, [253] introduced ChatCite, an LLM agent that extracts key elements from literature and generates summaries using a Reflective Incremental Mechanism for detailed summarization.
Retrieval-Augmented Generation for ATS: Retrieval-Augmented Generation (RAG) combines information retrieval with neural text generation, enabling models to generate text using both trained parameters and retrieved passages [274]. In ATS, RAG enhances summarization by supplementing external knowledge. [223] introduced Graph RAG, which uses a graph-based index to generate community summaries for comprehensive answers, improving performance on large datasets. Similarly, [254] proposed REDCODER to retrieve relevant code or summaries for code generation, and [249] developed a retrieve-then-summarize method using templates to summarize retrieved information. In the medical domain, [255] proposed a hybrid extractive-abstractive summarization method for processing large unstructured text, while [256] introduced LogicSumm and SummRAG to enhance LLM robustness in RAG-based scenarios. To reduce token consumption, [238] fine-tuned a T5 model to summarize and compress texts, and [239] designed extractive and abstractive compressors to improve task performance and selectively omit irrelevant content.
6.2.2 Fine-tuning LLMs for ATS
Given the massive number of parameters in LLMs, training the entire model requires substantial computational resources. To address this, parameter-efficient tuning techniques are often employed for fine-tuning on ATS datasets. These techniques involve either selectively training a subset of the LLM’s internal parameters or constructing and then fine-tuning external neural adapters to adapt the LLM to the ATS task.

### v3

@@ 6.1 Extractive Summarization
Extractive ATS based on supervised learning uses a classifier to categorize sentences as summaries or non-summaries. The process involves representing sentences, classifying them, and selecting those labeled as summaries. Classifiers from other tasks can be adapted for this purpose, with options including machine-learning-based (ML) and Pre-trained-based methods. While supervised approaches improve sentence representation and extraction accuracy using labeled data, they incur additional costs for annotation.
@@
Pre-trained-based Models: Pre-trained-based models offer rich textual representations that enhance sentence vector construction and improve summarization accuracy. [195] introduced SqueezeBERTSum, fine-tuned with a SqueezeBERT encoder variant, while [196] explored generative and discriminative techniques to fuse domain knowledge into knowledge adapters, integrating them into pre-trained models for efficient fine-tuning. However, training these models is computationally intensive, requiring significant hardware resources.
@@
6.2 Abstractive Summarization
Abstractive summarization models generate summaries by producing sentences distinct from the original text. This can be achieved through text structuring and combining or via generative models trained to predict the next token. Mainstream approaches of abstractive summarization are discussed as the following:
@@ 6.2 Abstractive Summarization
Prompt Engineering Template Engineering [15, 239, 240, 16, 241] Predictable output, easy to control structure. Limited flexibility, may not capture complex relationships.
@@
RAG - [215, 246, 241, 247, 248] Can use existing knowledge bases effectively. Reliability depends on retrieval accuracy; may introduce bias.
Fine-tuning Internal parameters [67, 249, 250, 251, 252, 253] Tailors model to specific tasks or domains. Destroying the original structure and capabilities of LLM.
External adapters [254, 255, 256, 257, 258, 259] Less risk of overfitting; can adapt quickly to new tasks. May not transfer improvements as effectively as full fine-tuning.
@@
6.3 Hybrid Summarization
7 Large Langue Models (LLMs) based Summarization Methods
@@
In ATS, LLMs have achieved results comparable to or exceeding human performance. Studies indicate that their summaries match human-crafted ones in quality and paraphrasing diversity [268, 269, 270, 271]. Moreover, when prompted with simple task descriptions, LLMs generate preferred and highly factual summaries, avoiding dataset-specific pitfalls [272]. This review bridges gaps in the literature by providing a comprehensive overview of LLMs in ATS.
LLM-based methods can be generally categorized as: 1) prompt engineering, 2)retrieval Augmented Generation, 3) fine-tuning, and 4) knowledge distillation. The advantages and limitations of these methods are summarized in Table 7.
Figure 2: Methodology Overview of Prompt Engineering for LLM-based ATS methods, including: (a) Template-Based Prompting aims to design structured prompt words to guide LLMs in generating effective summaries; (b) Chain-of-Thought (CoT) Prompting incorporates step-by-step reasoning to help generate summaries; (c) Multi-Agent Collaboration involves distinct LLMs for extraction, evaluation, and generation, enabling modular summarization workflows
7.1 Prompt Engineering for LLM-based ATS
@@
$S=LLM_{\theta}\left(G(p)+x\right).$ (3)
In the equation $S=LLM_{\theta}(G(p)+x)$, the function $G$ represents a refinement process that transforms an initial prompt $p$ into a more effective version using methods such as manual prompt engineering, retrieval of supplementary information, or automatic generation by another LLM. This refined prompt is then combined with the raw input text $x$ to guide the language model $LLM_{\theta}$, where $\theta$ denotes the model’s learned parameters. By enriching the input with a well-designed prompt, this approach helps the model better interpret the task and the context of the input, leading to improved quality in the generated summary $S$.
@@
The key advantage of prompt engineering is its efficiency, as it minimizes the need for extensive training and can operate effectively with a small set of examples [16]. We review three specific categories of approaches under the prompt engineering technique in ATS, including Template Engineering, Chain of Thought and Agent Interactions, as illustrated in Figure 2.
7.1.1 Template Engineering for ATS
Template engineering in text summarization employs structured linguistic scaffolds to guide LLMs in systematically extracting, condensing, and regenerating textual content while ensuring coherence, contextual relevance, and alignment with user-defined objectives. By integrating explicit instructions, contextual anchors, and output specifications, this approach leverages the semantic comprehension, discourse coherence, and domain adaptability of LLMs to produce summaries that preserve factual accuracy, emphasize key insights, and adhere to application-specific constraints, thereby enabling scalable and intent-aware summarization across diverse domains. The representative templates are as follows:
  • 1.
    $[Input]$ TL;DR: [15]
  • 2.
    Article: $[Input]$. Summarize the article in three sentences. Summary: [273]
  • 3.
In practice, [274] proposed a dual-template framework for news summarization by leveraging SPARQL queries on knowledge graphs, distinguishing between a simple template that concatenated subject-predicate-object triplets from document subgraphs and a complex template that incorporated subgraph-level knowledge inference and aggregation to mimic human reasoning. The design emphasized syntactic traversal of directional triplets for basic summarization and hierarchical task decomposition for advanced reasoning. To enhance the alignment of LLMs with the requirements of dialogue summarization tasks, [219] proposed Baichuan2-Sum, a role-oriented dialogue summarization model based on Baichuan2, which used a role-specific prompt template to address the limitations of small-scale models in capturing multi-role interactions. The design centered on assigning distinct instructions for different dialogue roles, enabling the model to learn role-dependent contextual patterns and generate targeted summaries.
Table 8: Prompt templates for mental health risks summarization extracted from [275].
As a mental health assistant, your task is to $[TASK]$ directly from the provided input text to highlight the mental health issues. For the $[TASK]$ task, consider the following aspects: 1. Emotions: Evaluate expressed emotions, from sadness to intense psychological pain, as they may influence the assigned risk level. 2. Cognitions: Explore the individual’s thoughts and perceptions about suicide, including the level and frequency of suicidal thoughts, intentions of suicide, and any existing plans. 3. Behavior and Motivation: Evaluate the user’s actions related to suicide, such as access to means and concrete plans. Consider their ability to handle difficult/stressful situations and the motivations behind their desire to die. 4. Interpersonal and Social Support: Investigate the individual’s social support or stable relationships, and understand their feelings toward significant others. 5. Mental Health-Related Issues: Consider psychiatric diagnoses associated with suicide such as schizophrenia, bipolar, anxiety, eating disorder, previous suicidal attempts, and others. 6. Additional Risk Factors: Consider other factors like socioeconomic and demographic factors, exposure to suicide behavior by others, chronic medical conditions, etc.
For healthcare community question answering, [276] introduced a novel perspective-specific answer summarization task, addressing the challenge of synthesizing diverse, multi-perspective responses into coherent summaries. To operationalize this, the authors designed a “structured prompt template” that explicitly defined perspective categories (e.g., suggestion), specified summary tone (e.g., advisory), and guided the model to generate summaries starting with predefined phrases (e.g., “It is suggested”), ensuring alignment with the target perspective. Furthermore, [275] investigated template prompting leveraging to detect and summarize evidence of suicidal ideation in social media, aiming to extract critical mental health-related information. The framework employed structured instruction templates that guided LLMs to focus on six dimensions shown in Table 8, including emotions, cognitions, behavior and motivation, interpersonal/social support, mental health issues, and additional risk factors. By providing explicit task-oriented instructions with contextualized examples, the study evaluated LLMs’ inherent capability to identify key text spans indicating mental health risks.
Table 9: Different prompt templates for ATS extracted from [277].
# Prompt Template
A Summarize in a single short paragraph the context below: $[document]$ The summary is: $[summary]$
B Summarize in a couple of sentences the document below: $[document]$ The summary is: $[summary]$
C Give me a short summary of the below: $[document]$ The summary is: $[summary]$
[…]
The effectiveness of these prompts is further amplified by the rapid advancement of LLMs, which possess enhanced capabilities in understanding nuanced semantic relationships, maintaining discourse coherence, and adapting to domain-specific terminology. As models grow in parameter count and training data diversity, the synergy between well-crafted prompts and model architecture ensures that summaries become increasingly precise, context-aware, and aligned with user intent, ultimately transforming text summarization into a more intelligent and scalable solution for both scholarly and industrial applications.
7.1.2 Chain of Thought for ATS
Chain of Thought (CoT) is a prompting strategy that structures the model’s reasoning process by decomposing complex summarization tasks into explicit, sequential inference steps. This approach guides LLMs to systematically analyze input text through intermediate reasoning, such as identifying key entities, discerning hierarchical relationships, and synthesizing core arguments, before generating a coherent summary. By embedding step-by-step logical derivations, CoT aligns the generative process of the model with human-like cognitive patterns, reducing hallucinations and enhancing factual consistency. This method is particularly effective in multi-step summarization scenarios, such as multi-document synthesis or domain-specific summarization tasks, where structured reasoning ensures outputs remain contextually grounded and aligned with user-defined objectives [280].
A typical paradigm of CoT involves prompting LLMs to first systematically extract key elements from the input text, followed by generating a summary based on these structured intermediate representations. Illustrative Cases from [232]:
@@ 7.1.2 Chain of Thought for ATS
    What are the important entities in this document? What are the important dates in this document? What events are happening in this document? What is the result of these events? Please answer the above questions:
@@
    Let’s integrate the above information and summarize the article:
Following this work, [281] adapted this paradigm to source code summarization by designing a code-specific CoT framework, the designed CoT included prompts:
  1. 1.
[…]
  2. 2.
    Get LLMs’ response to Instruction 1, i.e., Response 1.
  3. 3.
    Instruction 2: “Let’s integrate the above information and generate a short comment in one sentence for the function.”
  4. 4.
    Get LLMs’ response to Instruction 2, i.e., Response 2. Response 2 contains the comment generated by LLMs for the code snippet.
when asking Instruction 2, Instruction 1 and Response 1 are paired as history prompts and answers input into the LLM. By systematically addressing these questions through intermediate reasoning steps, the model generates summaries that preserve both syntactic accuracy and semantic coherence in source code contexts.
To balance conciseness and informativeness in summarization, [242] proposed a Chain of Density (CoD) prompting strategy to iteratively refine GPT-4 summaries. The CoT-inspired approach was motivated by the need to systematically address the tradeoff between entity coverage and summary readability, as overly sparse or dense outputs often fail to meet user expectations. The method involved first generating an entity-sparse summary via a standard prompt, then iteratively appending missing salient entities through a structured prompt sequence that maintained fixed output length while enhancing abstraction and fusion. The specific prompts are shown in the following case:
Example of Chain-of-Thought Prompting via “China of Density Prompting” extracted from [242] Article: ARTICLE{{ARTICLE}} You will generate increasingly concise, entity-dense summaries of the above Article. Repeat the following 2 steps 5 times. Step 1. Identify 1-3 informative Entities (“,” delimited) from the Article which are missing from the previously generated summary. step 2. Write a new, denser summary of identical length which covers every entity and detail from the previous summary plus the Missing Entities. Missing Entity is: - Relevant: to the main story. - Specific: descriptive yet concise(5 words or fewer). - Novel: not in the previous summary. - Faithful: present in the Article. - Anywhere: located anywhere in the Article. Guidelines: - The first summary should be long(4-5 sentences,  80 words) yet highly non-specific, containing little information beyond the entities marked as missing,Use overly verbose language and fillers(e.g.,“this article discusses”) to reach  80 words. - Make every word count: re-write the previous summary to improve flow and make space for additional entities. - Make space with fusion, compression, and removal of uninformative phrases like “the article discusses”. - The summaries should become highly dense and concise yet self-contained, e.g., easily understood without the Article. - Missing entities can appear anywhere in the new summary. - Never drop entities from the previous summary. If space cannot be made, add fewer new entities. Remember, use the exact same number of words for each summary. Answer in JSON, The JSON should be a list (length 5) of dictionaries whose keys are “Missing Entities” and “Denser Summary”
In addition, 3A-COT framework[282] was motivated by human-like reasoning patterns in MDS, leveraging CoT to decompose the task into three structured stages: Attend, Arrange, and Abstract. First, an Attend-prompt extracted key elements from each document to mitigate redundancy; second, an Arrange-prompt resolved contradictions and established logical connections between these elements; finally, an Abstract-prompt synthesized a coherent summary based on the arranged information. [227] addressed the challenges of clustering and abstractive summarization of online news comments by integrating a CoT strategy to enhance LLMs in generating coherent summaries for noisy, ambiguous user-generated content. The CoT approach was motivated by the need to systematize the model’s reasoning process when synthesizing dense, semantically diverse comment clusters, ensuring alignment with the structural and contextual nuances of real-world discussions.
The limitations of CoT in text summarization primarily stem from its computational demands and task-specific constraints: the generation of explicit reasoning steps significantly increases inference costs and delays response times compared to direct output methods. Additionally, CoT raises concerns about loyalty and transparency, as advanced models frequently obscure their reliance on external cues or shortcuts in reasoning explanations, creating a disconnect between internal computations and external justifications that undermines trust in model honesty[283].
7.1.3 Agent Interactions for ATS
Agent interactions frameworks first deconstruct the summarization task into sub-objectives, then iteratively refine outputs by leveraging internal memory and finally adjust strategies based on feedback from dynamic environments[284], enhance the autonomy and adaptability of LLMs by embedding task-specific reasoning, memory, and environmental interaction mechanisms.
Typically, [285] proposed a Multi-LLM summarization framework with two distinct strategies: centralized and decentralized. Both strategies involved k LLM agents generating diverse summaries, but differed in evaluation: centralized used a single central agent to select the best summary, while decentralized employed all k agents for evaluation. The implemented algorithm is shown in Algorithm 2. [286] introduced an innovative pipeline leveraging LLMs with over 100B parameters, such as GPT-3.5 and GPT-4, to function as synthetic feedback experts capable of generating high-quality edit feedback for enhancing factual consistency in clinical note summarization. The researchers implemented two distinct alignment algorithms to utilize this synthetic feedback, aiming to reduce hallucinations and align weaker, smaller-parameter LLMs (e.g., GPT-2 and Llama 2) with medical facts without requiring additional human annotations.
Algorithm 2 A case related to Agents Interactions extracted from [285].
Require: ordered set $S=\{S_{1},\ldots,S_{m}\}$ of summaries, set $\mathcal{M}=\{M_{1},\ldots,M_{k}\}$ of $k$ LLMs, a central agent $C\in\mathcal{M}$, max number of conversational rounds $t_{\max}$, initial summarization prompt $P$, evaluation prompt $P_{ec}$for centralized version
Ensure: summary $S^{*}$ of the text
[…]
2:for $i=1$ to $t_{\max}$ do $\triangleright$ conversation rounds
3:  for model $M_{j}\in\mathcal{M}$ do
4:   $S^{(i)}_{j}=M_{j}(P,S)$
5:  end for
6:  Let $\mathcal{S}_{i}=\{S^{(i)}_{1},S^{(i)}_{2},\ldots,S^{(i)}_{k}\}$
7:  $E^{(i)}=C(P_{ec},\mathcal{S}_{i})$
8:  $\mathbf{r}=\textsc{AggrResults}(E^{(i)})$
9:  $j\leftarrow argmax_{M_{j}\in\mathcal{M}}r_{j}$
10:  Set $S^{*}\leftarrow S^{(i)}_{j}$
11:  if Converged($\mathbf{r}$) then
12:   return $S^{*}$
13:  end if
14:  Set $P$ to prompt.
15:end for
@@ 7.1.3 Agent Interactions for ATS
In addition, [243] proposed a tri-agent pipeline including generator, instructor, and editor, to customize LLM-generated summaries to better meet user expectations. [99] used agents to address dialogue summarization challenges like lengthy inputs, content verification, limited annotated data, and evaluation, with GPT-3 serving as an offline data annotator to handle dataset scarcity and privacy constraints. [244] employed clustering-based agents to improve intent detection by abstracting key elements while removing irrelevant information. Similarly, [245] introduced ChatCite, an LLM agent that extracts key elements from literature and generates summaries using a Reflective Incremental Mechanism for detailed summarization.
Besides, summarization-oriented LLM can be integrated as an autonomous agent within a multi-agent system framework. For example, [287] proposed a novel agent framework for automating resume screening, motivated by the need to enhance efficiency and accuracy in recruitment processes through structured decision-making and scalable summarization of candidate profiles. The framework integrated automatic summarization as a core component, leveraging LLM agents to generate concise, graded summaries of resumes while aligning with domain-specific criteria such as skill matching and experience evaluation. [288] investigated the integration of summarization agents into interactive systems, motivated by the need to align generated outputs with user preferences through iterative refinement. The agent dynamically adjusted its output by retrieving and aggregating learned preferences from similar historical contexts, enabling personalized abstractions while preserving factual coherence in response to user-driven refinements. [289] proposed a planning framework for intelligent agents that utilized an automated summarization agent to extract and structure task-relevant knowledge from expert trajectories, enabling the synthesis of hierarchical state knowledge through comparative analysis and sequential abstraction to train a World Knowledge Model aligned with domain-specific task requirements.
In the context of summarization tasks, agents emerge as autonomous systems capable of integrating reasoning, planning, and tool utilization, distinguishing themselves from direct prompting and chain-of-thought (CoT) approaches. While direct prompting relies on static instructions to elicit model outputs and CoT enhances reasoning by structuring intermediate steps through human-designed prompts, agents operate dynamically, leveraging internal memory, environmental interaction, and adaptive decision-making to iteratively refine summaries. Their strengths lie in handling complex, multi-step workflows, such as cross-referencing external databases or synthesizing heterogeneous data sources, while maintaining contextual coherence and task-specific objectives. However, agents face challenges in interpretability, as their autonomous actions may obscure the rationale behind decisions. Additionally, compared to prompt-based methods, agents often demand higher computational resources and meticulous engineering to balance autonomy with precision, particularly in scenarios requiring strict adherence to factual accuracy or nuanced linguistic patterns. Thus, while agents excel in tasks demanding sustained reasoning and environmental engagement, their broader adoption hinges on addressing scalability, transparency, and robustness in real-world applications.
7.2 Retrieval Augmented Generation for ATS
Retrieval Augmented Generation (RAG) enhances the summarization generation process by incorporating externally retrieved documents to supplement the model’s internal knowledge [290], thereby improving the factual accuracy and coherence of summaries. Typically, as shown in Figure 3, this approach involves two stages: first, a retrieval mechanism such as BM25 or dense passage retrieval identifies semantically aligned segments from a pre-indexed knowledge base, and second, a sequence-to-sequence generator fuses these retrieved contexts with the input document to produce summaries that retain salient information while mitigating hallucinations. This dual-stage architecture ensures that the generated output is both linguistically fluent and grounded in verifiable external evidence.
Figure 3: Example from [291]. The RAG-based summarization methods first retrieve contextually relevant information from external knowledge bases using retrieval strategies, then generate abstractive summaries by integrating the retrieved content with the input document to ensure factual coherence and mitigate hallucinations.
Recent advancements in RAG have demonstrated diverse strategies for integrating external knowledge into generative models. [215] introduced Graph RAG, which constructs a graph-based index by extracting structured entities and relationships from text, followed by community detection algorithms to partition the graph into semantically coherent clusters. This framework retrieves community-level summaries derived from knowledge graphs, enabling global context-aware reasoning for complex queries on large-scale datasets. Similarly, [292] proposed Context-Aware Fine-Grained Graph RAG, extended traditional graph-based retrieval by introducing context-aware entity expansion, which broadened the coverage of retrieved entities through dynamic contextual relationships, and query-level fine-grained summarization, which incorporated detailed, query-specific interactions during generation to align summaries with latent user intent.
In highly specialized domains, where summarization tasks demand precise, domain-specific knowledge and up-to-date factual accuracy, Retrieval-Augmented Generation (RAG) plays a pivotal role by integrating external knowledge bases into the generation process. [293] aimed to address the challenges of navigating complex, high-volume experimental data in the Electron Ion Collider domain by developing an RAG-based summarization framework (RAGS4EIC), motivated by the need to streamline information access for researchers while preserving contextual accuracy and traceability. The RAGS4EIC design implemented a two-stage process: first, querying a vector database of structured experimental data to retrieve relevant context, followed by generating concise, citation-enriched summaries using an LLM, with additional flexibility provided by prompt-template-based instruction-tuning to align outputs with user-specific query intent. In code generation, [246] proposed REDCODER, employing dense retrieval to identify semantically similar code snippets or summaries from repositories, while [241] developed a retrieve-then-summarize pipeline that used sliding window segmentation and template-based abstraction to condense retrieved code fragments into task-specific outputs.
In the medical field, [247] designed a hybrid extractive-abstractive summarization approach, retrieving unstructured clinical texts (e.g., electronic health records or research papers) and integrating both direct evidence extraction and context-aware abstraction to generate domain-specific summaries. [248] introduced LogicSumm and SummRAG, which leverage retrieval from structured medical knowledge bases (e.g., clinical guidelines or pharmacological databases) to enhance logical consistency and factual accuracy in LLM-based reasoning tasks.[294] proposed the RAG-RLRC-LaySum framework to address the challenge of simplifying complex biomedical research for lay audiences, motivated by the need to bridge domain-specific jargon and enhance accessibility through structured knowledge integration. The design incorporated an RAG pipeline enhanced with a reranking mechanism, which aggregated diverse knowledge sources to ensure factual accuracy and contextual relevance in lay summaries, while a Reinforcement Learning for Readability Control component dynamically optimized linguistic simplicity and coherence during generation. [295] addressed the limitations of existing Electronic Health Records (EHR) models in capturing nuanced medical context by proposing REALM, an RAG-driven framework designed to integrate external knowledge and bridge gaps in clinical predictive tasks. REALM’s design combined LLM-based encoding of clinical notes with GRU-driven temporal modeling of EHR sequences, followed by a knowledge alignment module that extracted task-relevant entities from PrimeKG and fused them with multimodal data via an adaptive fusion network, ensuring semantic consistency and reducing hallucinations through contextual grounding in standardized medical knowledge. [296] designed a QA-based RAG framework to address the limitations of traditional neural models in EHR summarization, particularly the challenges of insufficient annotated training data and quadratic computational complexity due to long input lengths. The RAG design leveraged predefined questions identified by subject-matter experts to guide the extraction of targeted answers from EHRs, combining semantic search for relevant content retrieval with LLM-based generation to ensure factual consistency and reduce hallucinations, while avoiding the need for extensive training.
Retrieval-augmented generation (RAG) for ATS integrates external knowledge retrieval with generative models to enhance factual accuracy and contextual relevance. By dynamically accessing external databases or domain-specific corpora, RAG mitigates the limitations of standalone language models, which may generate hallucinated or outdated information due to fixed training data. This approach ensures summaries are factually grounded, particularly in specialized domains such as legal or scientific texts, where access to authoritative references improves coherence and reliability. Furthermore, RAG introduces interpretability by anchoring summaries to retrievable sources, enabling traceable reasoning. However, its effectiveness is constrained by the quality and coverage of the retrieval system; incomplete or noisy knowledge bases can propagate errors into summaries. Additionally, the dual-stage pipeline—retrieval followed by generation—introduces computational overhead and latency, limiting real-time deployment in resource-constrained settings. The interplay between retrieved content and generative processes may also result in fragmented outputs if contextual integration is suboptimal, undermining fluency. While RAG addresses critical gaps in model knowledge and bias, its practical utility hinges on robust knowledge curation, efficient retrieval mechanisms, and seamless alignment of retrieved information with the summarization objective.
7.3 Fine-tuning LLMs for ATS
Fine-tuning is pivotal for tailoring LLMs to specialized domains such as legal systems, medical diagnostics, and scientific research. This process significantly enhances their ability to grasp domain-specific knowledge and adapt to field-specific linguistic conventions. As shown in Figure 4, the approaches typically fall into two categories: (1) Internal parameter‑efficient fine‑tuning, which strategically activates and updates only a subset of the LLM’s internal parameters while preserving its pre-trained core capabilities; and (2) External adapters fine-tuning, where compact neural modules are inserted between transformer layers to learn domain-specific transformations[267]. By focusing computational resources on these targeted adaptation mechanisms, fine-tuning enables LLMs to effectively capture ATS-relevant features such as semantic hierarchies, discourse structures, and task-specific terminology, while maintaining general language understanding capabilities.
Figure 4: Example from [267]. Fine-tuning LLMs for ATS are primarily categorized into internal fine-tuning approaches that selectively update a subset of pre-trained parameters to retain foundational linguistic knowledge and external adapters fine-tuning that employ compact neural modules inserted between transformer layers to encode domain-specific transformations.
7.3.1 Internal efficient-parameters fine-tuning